# Heterogeneous model for superdiffusive movement of dense-core vesicles in C. elegans

### v1

@@ Acknowledgements
A.G. generated specific worm strains, imaging neurons with a spinning-disk confocal microscope, analysed DCV movement using KymoButler and custom Python code, segmented and calculated displacements, fitted a beta-binomial distribution to the data, prepared figures, and co-authoring the text of the paper. N.K. contributed to formal analysis, validation, investigation, and writing (review and editing). V.A. contributed to conceptualization, resources, data curation, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing). S.F. contributed to conceptualization, formal analysis, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing).

### v2

@@ Acknowledgements
A.G. generated specific worm strains, imaged neurons with a spinning-disk confocal microscope, analysed DCV movement using KymoButler and custom Python code, segmented and calculated displacements, fitted a beta-binomial distribution to the data, prepared figures, and co-authoring the text of the paper. N.K. contributed to formal analysis, validation, investigation, and writing (review and editing). V.A. contributed to conceptualization, resources, data curation, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing). S.F. contributed to conceptualization, formal analysis, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing).
@@
Data availability: The data supporting this study’s findings are available upon request from the corresponding author.